# AdaptiveK Sparse Autoencoders: Dynamic Sparsity Allocation for Interpretable LLM Representations

### v1

@@ Abstract
Understanding the internal representations of large language models (LLMs) remains a central challenge for interpretability research. Sparse autoencoders (SAEs) offer a promising solution by decomposing activations into interpretable features, but existing approaches rely on fixed sparsity constraints that fail to account for input complexity. We propose Adaptive Top K Sparse Autoencoders (AdaptiveK), a novel framework that dynamically adjusts sparsity levels based on the semantic complexity of each input. Leveraging linear probes, we demonstrate that context complexity is linearly encoded in LLM representations, and we use this signal to guide feature allocation during training. Experiments across three language models (Pythia-70M, Pythia-160M, and Gemma-2-2B) demonstrate that this complexity-driven adaptation significantly outperforms fixed-sparsity approaches on reconstruction fidelity, explained variance, and cosine similarity metrics while eliminating the computational burden of extensive hyperparameter tuning.
@@ 1 Introduction
How can we define semantic simplicity or complexity within LLM representation spaces? Peters et al. (2018) demonstrated that probes can map LLM intermediate representations to semantic and syntactic information. Similarly, higher-level concepts such as political perspective (Kim et al., 2025), sentiment (Tigges et al., 2023), and spatiotemporal information (Gurnee and Tegmark, 2023) have been shown to be linearly represented in activation spaces. Based on these observations, we hypothesize that the internal representations formed by large language models during text processing naturally encode multidimensional properties of text, including its complexity.
Our study first establishes that text complexity is linearly encoded in language model representations. We score contexts using GPT-4.1-mini API (OpenAI, 2024) across six semantic dimensions to create aggregate complexity scores, then train linear regression probes on model activations to predict these scores. Experiments with three different scale LLMs demonstrate high correlation coefficient, confirming our hypothesis that LLMs naturally encode text complexity in their representation spaces. Analysis reveals that texts of varying complexity require proportional representational capacity, with complex inputs necessitating more active features for accurate encoding. This key insight suggests that adaptive sparsity mechanisms could significantly improve autoencoder efficiency.
Based on our complexity prediction capabilities, we develop the Adaptive Top K Sparse Autoencoder (AdaptiveK SAE), which to our knowledge is the first work to solve computational efficiency bottlenecks in sparse autoencoder training while maintaining feature quality. Our approach first quantifies the inherent complexity of each context using a linear probe trained on multi-dimensional complexity annotations. This complexity score then determines an appropriate sparsity level, allowing more features to activate for complex inputs while maintaining high sparsity for simpler ones. This complexity-driven adaptation achieves an improved balance between reconstruction quality, sparsity, and interpretability. Experiments demonstrate that our framework outperforms fixed-sparsity approaches across multiple model scales. Our main contributions can be summarized as follows:
@@
  • Experiments across three LLMs show that our SAE consistently outperforms fixed-sparsity baselines on reconstruction fidelity, explained variance, cosine similarity and other metrics.
@@ 2.2 Our Motivation
Despite advances in SAE scalability and architecture, current approaches face a fundamental limitation: they apply uniform sparsity constraints regardless of input complexity. This “one-size-fits-all” approach creates inefficiencies across the representation space. Whether employing fixed activation methods such as TopK (Gao et al., 2024) or regularization techniques like Gated SAEs (Rajamanoharan et al., 2024a) existing architectures cannot adapt to the varying complexity demands of different inputs. As shown in Figure 1 conventional SAE methods with fixed K maintain constant activation (e.g., 80 features across all complexities), while our AdaptiveK dynamically scales from 103 to 394 features based on input complexity. This inefficiency becomes more pronounced at scale, determining optimal sparsity parameters necessitates extensive hyperparameter optimization to balance reconstruction fidelity against sparsity constraints.
@@
Figure 1: Two samples were selected from each complexity level (simplest=0, moderate=5.5, and most complex=9.6) from test set. In TopK SAE, feature activation strictly follows the k value between 20-320, but often falls below the threshold when k=640. For Pythia-160M, fixed TopK SAEs (blue) maintain constant activation, while AdaptiveK (red) dynamically scales with context complexity.
@@ 3 The Proposed AdaptiveK Sparse Autoencoder
In this work, we propose AdaptiveK Sparse Autoencoder that dynamically adjusts sparsity constraints according to input complexity. Our AdaptiveK Sparse Autoencoder addresses the fundamental inefficiency of uniform sparsity constraints by allocating representational capacity proportionally to content complexity (see Figure 2). Specifically, the AdaptiveK SAE consists of two primary components: (1) a linear probe that predicts input complexity (Section 3.1), and (2) the sparse autoencoder with AdaptiveK activation (Section 3.2). Additionally, we introduce our three-phase scheme for effective training of our proposed AdaptiveK SAE (Section 3.3).
Figure 2: Overall pipeline of the AdaptiveK SAE framework. The text input is fed into a LLM to extract internal activations, which are then passed through both a linear probe that predicts text complexity and a SAE for decomposition. During the training process of the SAE, the linear probe’s complexity score dynamically determines the number of features to activate, allowing more features for complex inputs and fewer for simple ones.
@@ 3.1.1 Linear Probe for Complexity Prediction
Our dataset comprises contexts from pile-uncopyrighted (Gao et al., 2020), each containing 1024 tokens. Complexity is quantified on a scale from 0 to 10, based on a six-dimensional evaluation (lexical complexity, syntactic complexity, conceptual density, domain specificity, logical structure and logical structure) by GPT-4.1-mini (scoring prompts detailed in Appendix A), yielding target labels $y_{i}$. These complexity scores are floating-point numbers with one decimal place. For each context, we extract the hidden state activations of the last token from selected layers of the auto-regressive transformer language models Pythia-70M, Pythia-160M (Biderman et al., 2023) and Gemma-2-2B (Team et al., 2024), which encapsulate contextual information. The input data for each context $i$ is represented as a pair $x_{i},y_{i}]$, where ${x_{i}}\in\mathbb{R}^{d_{\text{model}}}$ is the vector of hidden state activations and $y_{i}$$\in\mathbb{R}$ is the corresponding complexity score.
Following Gurnee and Tegmark (2023), we employ ridge regression to mitigate overfitting and multicollinearity issues common with high-dimensional activation vectors (Kim et al., 2025). The objective is to find the weight vector $w\in\mathbb{R}^{d_{\text{model}}}$ and bias term $b\in\mathbb{R}$ that minimize the $L_{2}$-regularized squared loss:
@@
where $n$ is the number of training contexts, and $\lambda$ is the regularization hyperparameter. The activation matrix $A\in\mathbb{R}^{n\times d_{\text{model}}}$ is constructed by concatenating these feature vectors, with the target vector defined as $y\in\mathbb{R}^{n}$. With implicit bias handling (e.g., by centering the data or adding a feature column of ones to $A$), the optimal weight vector $\hat{w}$ is given by the closed-form solution (Belinkov, 2022):
@@
(b) Pythia-160M
(c) Gemma-2-2B
Figure 3: Visualization of linear probe performance across three LLM scales. Points represent test contexts, with redder areas indicating higher sample density. The red line depicts predicted complexity trends. The red line depicts predicted complexity trends. Most samples fall within prediction intervals, confirming the linear probe’s effectiveness. Spearman Correlation and RMSE values (upper left) demonstrate improved prediction accuracy with increasing model scale.
@@ 3.1.2 Evaluating the Linear Representation Hypothesis
As will be mentioned in Section 5.2 , a growing number of high-level features have been demonstrated to exist linearly in LLM representation spaces. However, context complexity is indeed a multifaceted construct: it spans multiple linguistic dimensions, including lexical, syntactic complexity, etc., so it differs fundamentally from the single-attribute features investigated in previous work.
@@
As shown in Table 1, linear model is comparable to theoretically more expressive nonlinear models like XGBoost across all three metrics: Pearson correlation coefficient, Spearman correlation coefficient, and RMSE. These findings extend the applicability of the linear representation hypothesis to multifaceted features such as context complexity. Figure 3 visualizes our probe’s performance on test data across three models. Each point represents one context from the test set, showing the predicted versus true complexity score. With most points falling within the prediction interval (±1.96 $\times$ RMSE), this performance confirms that context complexity is largely encoded linearly in representation space.
@@ 3.2 AdaptiveK Sparse Autoencoder Architecture
For an input activation vector $x\in\mathbb{R}^{d}$, we first use the linear probe described in Section 3.1 to compute a complexity score $c$: $c=\hat{w}^{T}x+\hat{b}$, where $\hat{w}\in\mathbb{R}^{d}$ and $\hat{b}\in\mathbb{R}$ are trained from the ridge regression. This complexity score is then mapped to a sparsity level $k_{\mathrm{adp}}$ through a sigmoid-based transformation:
$k_{\mathrm{adp}}=k_{\min}+\frac{1}{1+e^{-s\cdot\left(\frac{c-c_{\min}}{c_{\max}-c_{\min}}-0.5\right)}}\cdot\left(k_{\max}-k_{\min}\right),$ (7)
@@ 12  Update all parameters using $L_{joint}=L_{SAE}+\gamma(L_{probe}+\delta L_{deviation})$
Our implementation uses a AdaptiveKBuffer that extracts last-token representations from contexts (see Section 3.1.1 for details), reducing memory usage while tracking complexity scores for balanced training. Adam optimizer (Kingma and Ba, 2014) is applied with learning rate 1e-3, warm-up over 15 steps, and linear decay starting at 70% of training. More SAE training details will be presented in Appendix B.
@@ 4.1 Experimental Settings
We train SAEs on three language models of increasing scale: Pythia-70M, Pythia-160M, and Gemma-2-2B (Biderman et al., 2023; Team et al., 2024), with hidden state dimensions of 512, 768, and 2048 respectively. All trained SAEs have a dictionary size of 16384, i.e., 16k latents. Unlike previous token-level approaches, we operate at the context level, using both training and evaluation data from pile-uncopyrighted (Gao et al., 2020). Our training dataset comprises 250,000 contexts, while our test set contains 10,000 contexts, each consisting of 2048 tokens. Following studies on layer-wise functionality (Gao et al., 2024), we extract activations from layer 3 of Pythia models and layer 12 of Gemma-2-2B. We train on the last token representation of each context, which efficiently captures the accumulated contextual information for complexity-driven sparsity adaptation.
For our experimental evaluation, we compare our AdaptiveK SAE against seven SAE baselines, including: (1) ReLU SAEs (Bricken et al., 2023), the foundational approach using ReLU activation with L1 penalty; (2) refined ReLU_new SAEs (Anthropic Interpretability Team, 2024); (3) TopK SAEs (Gao et al., 2024), which select the K highest activations; (4) BatchTopK SAEs (Bussmann et al., 2024a), extending TopK across batches; (5) Gated SAEs (Rajamanoharan et al., 2024a), which decouple feature selection from magnitude estimation; (6) P-anneal SAEs (Karvonen et al., 2024), using annealing $L_{p}$ norm penalties; and (7) Matryoshka SAEs (Bussmann et al., 2024b), which train nested dictionaries with increasing capacity.
In addition to the primary metrics discussed in the main text, we utilized a range of metrics from SAEBench (Karvonen et al., 2025), such as ‘Spurious Correlation Removal’ among others, to evaluate AdaptiveK SAE against baseline SAEs. Due to limited space in the main document, further details regarding these evaluations are provided in Appendix C.2. Furthermore, Appendix C.1 explores how the behavior of AdaptiveK SAE differ when it is applied to distinct layers within a single LLM.
@@ 4.2 Relationship Between Complexity and k-Values
We plotted the relationship between the true complexity and k-value selection on the test set, calculated the average k-value for each complexity interval, and marked the fixed k-values that TopK would select. As shown in Figure 4, there is an approximately linear relationship between text complexity and allocated k-values, which becomes increasingly evident as model scale increases.
On one hand, examining Figure 3, we observe that both predicted complexity and activated feature count increase with sample complexity. This validates our approach’s effectiveness: complex texts are accurately assessed and mapped to larger k values, activating more features. The results demonstrate the coherence of our complexity-aware sparse autoencoder framework.
On the other hand, considering Figure 5(b), 6(b) and 7(b), we observe that while reconstruction errors of TopK SAE gradually decrease as k-values increase, they remain consistently higher than those of AdaptiveK SAE. This demonstrates that AdaptiveK SAE intelligently allocates computational resources by assigning fewer features to simpler texts and more features to complex ones, ultimately achieving superior performance.
(a) Pythia-70M
(b) Pythia-160M
(c) Gemma-2-2B
Figure 4: Visualization of Dynamic Feature Allocation by Text Complexity showing the relationship between complexity scores and allocated feature counts (K values). Average K values per complexity interval (connected by red lines) demonstrate that more complex texts receive higher K allocations. This relationship becomes increasingly linear as LLM scale grows. Horizontal lines indicate fixed Standard TopK baselines with K values noted on the right.

### v2

@@ Abstract
Understanding the internal representations of large language models (LLMs) remains a central challenge for interpretability research. Sparse autoencoders (SAEs) offer a promising solution by decomposing activations into interpretable features, but existing approaches rely on fixed sparsity constraints that fail to account for input complexity. We propose AdaptiveK SAE (Adaptive Top K Sparse Autoencoders), a novel framework that dynamically adjusts sparsity levels based on the semantic complexity of each input. Leveraging linear probes, we demonstrate that context complexity is linearly encoded in LLM representations, and we use this signal to guide feature allocation during training. Experiments across ten language models (from 70M to 14B parameters) demonstrate that this complexity-driven adaptation significantly outperforms fixed-sparsity approaches on reconstruction fidelity, explained variance, cosine similarity and interpretability metrics while eliminating the computational burden of extensive hyperparameter tuning.
@@ 1 Introduction
How can we define semantic simplicity or complexity within LLM representation spaces? Peters et al. (2018) demonstrated that probes can map LLM intermediate representations to semantic and syntactic information. Similarly, higher-level concepts such as political perspective (Kim et al., 2025), sentiment (Tigges et al., 2023), and spatiotemporal information (Gurnee & Tegmark, 2023) have been shown to be linearly represented in activation spaces. Based on these observations, we hypothesize that the internal representations formed by large language models during text processing naturally encode multidimensional properties of text, including its complexity.
Our study first establishes that text complexity is linearly encoded in language model representations. We score contexts using GPT-4.1-mini API (OpenAI, 2024) across six semantic dimensions to create aggregate complexity scores, then train linear regression probes on model activations to predict these scores. Experiments with eight different scale LLMs demonstrate high correlation coefficient, confirming our hypothesis that LLMs naturally encode text complexity in their representation spaces. Analysis reveals that texts of varying complexity require proportional representational capacity, with complex inputs necessitating more active features for accurate encoding. This key insight suggests that adaptive sparsity mechanisms could significantly improve autoencoder efficiency.
Based on our complexity prediction capabilities, we develop the Adaptive Top K Sparse Autoencoder (AdaptiveK SAE), which to our knowledge is the first work to solve computational efficiency bottlenecks in sparse autoencoder training while maintaining feature quality. Our approach quantifies context complexity using a linear probe trained on multi-dimensional complexity annotations. This score determines an appropriate sparsity level, activating more features for complex inputs while maintaining high sparsity for simpler ones. This complexity-driven adaptation better balances reconstruction quality, sparsity, and interpretability. Experiments demonstrate our framework outperforms fixed-sparsity approaches across multiple model scales. Code is available at: https://anonymous.4open.science/r/adaptiveK-5258. Our main contributions can be summarized as:
@@
  • Experiments across ten LLMs show that our SAE consistently outperforms fixed-sparsity baselines on reconstruction fidelity, explained variance, cosine similarity and other metrics.
@@ 2.2 Our Motivation
Figure 1: Two samples were selected from each complexity level (simplest=0, moderate=5.5, and most complex=9.6) from test set. In TopK SAE, feature activation strictly follows the k value between 20-320, but often falls below the threshold when k=640. For Pythia-160M, fixed TopK SAEs (blue) maintain constant activation, while AdaptiveK (red) dynamically scales with text complexity.
Despite advances in SAE scalability, current approaches face a limitation: they apply uniform sparsity constraints regardless of input complexity. This “one-size-fits-all” approach creates inefficiencies across the representation space. Whether employing fixed activation methods such as TopK (Gao et al., 2024) or regularization techniques like Gated SAEs (Rajamanoharan et al., 2024a) existing architectures cannot adapt to the varying complexity demands of different inputs. As shown in Figure 1 conventional SAE methods with fixed K maintain constant activation (e.g., 80 features across all complexities), while our AdaptiveK dynamically scales from 103 to 394 features based on input complexity. This inefficiency becomes more pronounced at scale, determining optimal sparsity parameters necessitates extensive hyperparameter optimization to balance reconstruction fidelity against sparsity constraints.
@@ 3 The Proposed AdaptiveK Sparse Autoencoder
In this work, we propose AdaptiveK Sparse Autoencoder that dynamically adjusts sparsity to input complexity. By allocating representational capacity in proportion to content complexity, AdaptiveK addresses the core inefficiency of uniform sparsity (Figure 2). The architecture comprises two components: (1) a linear probe that predicts input complexity (Section 3.1); and (2) an SAE with AdaptiveK activation (Section 3.2). We also present a three-phase training procedure that stabilizes and improves learning (Section 3.3).
Figure 2: Overall pipeline of the AdaptiveK SAE. Input text is fed into a LLM to extract internal activations, which are then passed through both a linear probe that predicts text complexity and a SAE for decomposition. During training, the linear probe’s complexity score dynamically determines the number of features to activate, allowing more features for complex inputs and fewer for simple ones.
@@ 3.1.1 Linear Probe for Complexity Prediction
Our dataset comprises contexts from pile-uncopyrighted (Gao et al., 2020), each containing 1024 tokens. Complexity is quantified on a scale from 0 to 10, based on a six-dimensional evaluation (lexical complexity, syntactic complexity, conceptual density, domain specificity, logical structure and logical structure) by GPT-4.1-mini (scoring prompts detailed in Appendix B), yielding target labels $y_{i}$. These complexity scores are floating-point numbers with one decimal place. For each context, we extract the hidden state activations of the last token from selected layers of the auto-regressive transformer language models Pythia (70M, 160M) (Biderman et al., 2023), Gemma-2 (2B, 9B) (Team et al., 2024), Llama-3.1 (8B) (Grattafiori et al., 2024), Qwen-3 (8B, 14B) (Yang et al., 2025) and Phi-4 (14B) (Abdin et al., 2024), which encapsulate contextual information. The input data for each context $i$ is represented as a pair $x_{i},y_{i}]$, where ${x_{i}}\in\mathbb{R}^{d_{\text{model}}}$ is the vector of hidden state activations and $y_{i}$$\in\mathbb{R}$ is the corresponding complexity score.
Following Gurnee & Tegmark (2023), we employ ridge regression to mitigate overfitting and multicollinearity issues common with high-dimensional activation vectors (Kim et al., 2025). The objective is to find the weight vector $w\in\mathbb{R}^{d_{\text{model}}}$ and bias term $b\in\mathbb{R}$ that minimize the $L_{2}$-regularized squared loss:
@@
where $n$ is the number of training contexts, and $\lambda$ is the regularization hyperparameter. The activation matrix $A\in\mathbb{R}^{n\times d_{\text{model}}}$ is constructed by concatenating these feature vectors, with the target defined as $y\in\mathbb{R}^{n}$. With implicit bias handling (e.g., by centering data or adding a feature column of ones to $A$), the optimal weight vector $\hat{w}$ is given by the closed-form solution (Belinkov, 2022):
@@
(b) Gemma-2-2B
(c) Qwen-3-8B
(d) Phi-14B
Figure 3: Visualization of linear probe performance across different LLM scales. Points represent test contexts, with redder areas indicating higher sample density. The red line depicts predicted complexity trends. Most samples fall within prediction intervals, confirming the linear probe’s effectiveness. Spearman Correlation and RMSE values (upper left) demonstrate improved prediction accuracy with increasing model scale. More LLM results are in Figure 10.
@@ 3.1.2 Evaluating the Linear Representation Hypothesis
As will be mentioned in Appendix A.2 , many high-level features have been demonstrated to exist linearly in LLM representation spaces. However, context complexity is indeed multifaceted, spanning lexical, syntactic, and other linguistic dimensions, and thus differs from the single-attribute features studied in prior work.
@@
As shown in Table 1, linear model is comparable to theoretically more expressive nonlinear models like XGBoost across all three metrics: Pearson and Spearman correlations and RMSE. These findings extend the applicability of the linear representation hypothesis to multifaceted features such as context complexity. Figure 3 visualizes our probe’s performance on test data across eight models. Each point represents one context from the test set, showing the predicted versus true complexity score. With most points falling within the prediction interval (±1.96 $\times$ RMSE), this performance confirms that context complexity is largely encoded linearly in representation space.
@@ 3.2 AdaptiveK Sparse Autoencoder Architecture
For an input activation vector $x\in\mathbb{R}^{d}$, we compute a complexity score using the linear probe from Section 3.1: $c=\hat{w}^{T}x+\hat{b}$, where $\hat{w}\in\mathbb{R}^{d}$ and $\hat{b}\in\mathbb{R}$ are trained from the ridge regression. This score is then mapped to a sparsity level $k_{\mathrm{adp}}$ through a sigmoid-based transformation:
$k_{\mathrm{adp}}=k_{\min}+\frac{1}{1+e^{-s\!\left(\frac{c-c_{\min}}{c_{\max}-c_{\min}}-0.5\right)}}\,(k_{\max}-k_{\min}).$ (7)
@@
(a) Pythia-70M
(b) Gemma-2-2B
(c) Qwen-3-8B
(d) Phi-14B
Figure 4: Visualization of Dynamic Feature Allocation by Text Complexity showing the relationship between complexity scores and allocated feature counts (K values). Average K values per complexity interval (connected by red lines) demonstrate that complex texts receive higher K allocations, with this relationship becoming increasingly linear as LLM scale grows. Horizontal lines indicate fixed Standard TopK baselines with K values on the right. More LLM results are in Figure 11.
@@ 12   Update all parameters using $L_{joint}=L_{SAE}+\gamma(L_{probe}+\delta L_{deviation})$
Our implementation uses a AdaptiveKBuffer that extracts last-token representations from contexts (see Section 3.1.1 for details), reducing memory usage while tracking complexity scores for balanced training. Adam optimizer (Kingma & Ba, 2014) is applied with learning rate 1e-3, warm-up over 15 steps, and linear decay starting at 70% of training. More SAE training details are in Appendix C.
@@ 4 Experiments
  • RQ4: How is LLM interpretability defined, and how can the interpretability of AdaptiveK SAE be measured at a human-understandable level?
@@ 4.1 Experimental Settings
We train SAEs on 8 language models of increasing scale: Pythia (70M, 160M), Gemma-2 (2B, 9B), Llama-3.1 (8B), Qwen-3 (8B, 14B) and Phi-4 (14B), with hidden dimensions from 512 to 5120. All trained SAEs have a dictionary size of 16384, i.e., 16k latents. Unlike token-level approaches, we operate at context level (though we demonstrate in Section H that token-level evaluation is also applicable) using pile-uncopyrighted data (Gao et al., 2020): 250,000 training contexts and 10,000 test contexts, each with 2048 tokens. We train on the last token representation of each context, which efficiently captures accumulated contextual information for complexity-driven sparsity adaptation.
For experimental evaluation, we compare our AdaptiveK SAE against 7 SAE baselines: (1) ReLU SAEs (Bricken et al., 2023), using ReLU activation with L1 penalty; (2) refined ReLU_new SAEs (Anthropic Interpretability Team, 2024); (3) TopK SAEs (Gao et al., 2024), which select the K highest activations; (4) BatchTopK SAEs (Bussmann et al., 2024a), extending TopK across batches; (5) Gated SAEs (Rajamanoharan et al., 2024a), decoupling feature selection from magnitude estimation; (6) P-anneal SAEs (Karvonen et al., 2024), using annealing $L_{p}$ norm penalties; and (7) Matryoshka SAEs (Bussmann et al., 2024b), training nested dictionaries with increasing capacity.
Beyond the primary metrics, we evaluated AdaptiveK using SAEBench metrics (Karvonen et al., 2025) (Appendix F.3). Additional analyses include layer-wise performance (Appendix F.1), extensions to encoder-only and encoder-decoder models (Appendix F.2), hyperparameter analysis (Appendix F.4), and training efficiency analysis (Appendix G).
(a) Pythia-70M
(b) Gemma-2-2B
(c) Qwen-3-8B
(d) Phi-14B
Figure 5: L2 Loss pareto frontier results. More LLM results are in Figure 12.
(a) Pythia-70M
(b) Gemma-2-2B
(c) Qwen-3-8B
(d) Phi-14B
Figure 6: Unexplained Variance pareto frontier results. More LLM results are in Figure 13.
(a) Pythia-70M
(b) Gemma-2-2B
(c) Qwen-3-8B
(d) Phi-14B
Figure 7: Cosine Similarity pareto frontier results. More LLM results are in Figure 14.
@@ 4.2 Relationship Between Complexity and k-Values
We plotted the relationship between true complexity and k-value selection on the test set, calculated the average k-value for each complexity interval, and marked the fixed k-values that TopK would select. As shown in Figure 4, there is an approximately linear relationship between text complexity and allocated k-values, which becomes increasingly evident as model scale increases.
On one hand, Figure 3 reveals that both predicted complexity and activated feature count increase with sample complexity, validating that complex texts correctly receive larger k values and activate more features. On the other hand, Figure 5 indicate that that AdaptiveK consistently achieves lower reconstruction errors than TopK SAE across all k-values, demonstrating the benefit of adaptive resource allocation by using fewer features for simple texts and more for complex ones. Additional probe evaluations using PCA and layer-wise analysis are in Appendices E.1 and E.2.